Required information (The following information applies to the questions displayed below.) A food manufacturer reports the fo1. Compute return on investment. 2. Compute profit margin. 3. Compute investment turnover for the year. Complete this questioRequired 1 Required 2 Required 3 Compute profit margin. (Enter your answers in millions.) Profit Margin 1 Choose Denominator:Required 1 Required 2 Required 3 Compute investment turnover for the year. (Enter your answers in millions.) Investment TurnoAssume that each of the companys divisions has a required rate of return of 8%. Compute residual income for each division (E
Answer #1
1
Return on Investment
Choose Numerator: / Choose Denominator = Return on Investment
Investment Center Operating income / Average invested assets = Return on Investment
Beverage 358 / 2641 = 13.56%
Cheese 643 / 4441 = 14.48%
2
Profit Margin
Choose Numerator: / Choose Denominator = Profit margin
Investment Center Operating income / Sales = Profit margin
Beverage 358 / 2690 = 13.31%
Cheese 643 / 3934 = 16.34%
3
Investment Turnover
Choose Numerator: / Choose Denominator = Investment Turnover
Investment Center Sales / Average invested assets = Investment Turnover
Beverage 2690 / 2641 = 1.02
Cheese 3934 / 4441 = 0.89
4
($ millions) Beverage Cheese
Average Assets 2641 4441
Targeted return 8% 8%
Target income 211 355
Residual Income Beverage Cheese
Operating income 358 643
Less: Target income 211 355
Residual Income 147 288
Note: For 4, Target income has been rounded off. The answer will slightly differ if not rounded off